### v2

# Fundamental Oscillations of Massive Boson Stars and Distinguishability

Abstract
Massive Boson Stars are self-gravitating configurations of self-interacting scalar fields. The equation of state of massive boson stars and their masses, radii, modeled by a self-interacting scalar field with potential of the form $V(\phi)=\frac{1}{2}m^{2}|\phi|^{2}+\frac{1}{4}\lambda|\phi|^{4}$ are known to follow scaling relations. The non-radial fundamental oscillations of such massive BSs have been studied only for a few select model parameters so far. In this work, we demonstrate for the first time that the $f$-mode characteristics also follow a scaling in the strong interaction limit ($\lambda\gg m^{2}/M_{Pl}^{2}$). This opens up the outstanding prospect of studying the $f$-modes of massive BSs throughout the scalar DM parameter space. We study the implications of this finding by carrying out a detailed study of massive BS $f$-modes in a separate work. Here, having introduced this scaling, we use it to compare boson star oscillations with the neutron star and black hole quasinormal modes, thus providing a smoking gun for the distinguishability of BSs using gravitational waves.
Boson Stars — Exotic Compact Objects — Oscillations — Gravitational Waves
I Introduction
Dark matter (DM) and its nature are among the most intriguing open questions of physics. DM forms $\sim 25\%$ of the energy budget of our universe. However, its particle nature has remained elusive. DM candidate masses span 90 orders of magnitude from ultralight fuzzy DM to massive compact halo objects (MaCHOs), and primordial black holes (PBHs) [8]. Historically, the most popular candidate for DM has been Weakly Interacting Massive Particles (WIMPs). However, years of experimental effort in the search for WIMPs have been fruitless so far.
On the other hand, the discovery of Higgs Boson has been one of the most important discoveries, theoretically motivated by shortcomings of the standard model (SM) of particle physics. This discovery also established the existence of scalar (spin-0) bosons. Another problem in particle physics, the well-known strong CP violation, has resulted in the prediction of a new scalar particle, axion, which makes a natural candidate for DM (See [9, 15] for a recent review on Axion DM). As a consequence, DM particles such as light bosonic particles like axions, axion-like particles (ALPs), fuzzy DM, etc., have attracted the spotlight lately.
Boson stars (BSs) are stellar configurations of bosonic particles (see [51, 30] for updated review on BSs). These can form from primordial scalar fluctuations, gravothermal collapse in an early matter-dominated era [39], or via a process called gravitational cooling [41]. They can also be formed via trapping of DM in objects like stars and planets [19, 23, 22, 24]. BS is one of the possible exotic compact objects (ECOs) that can mimic other compact objects like black holes (BHs) and neutron stars (NSs). BSs, if they exist, would also contribute to the total DM budget (see [13] for a review on ECOs) of the universe. Their existence would be a key to the DM problem. Thus, observation of BSs can revolutionize our understanding of the universe.
Conventional electromagnetic telescopes can only probe BSs if they are accreting matter [33]. This method is inaccessible in the case of non-accreting BSs. However, the successful detection of gravitational waves [1] has opened up a new window to probe the dark sectors of the universe. Further, the planning of the next-generation detectors covering a wide frequency range makes the study of all possible GW sources of BSs very crucial and timely.
There have been considerable efforts to find observational signatures that will be able to distinguish BSs from BHs.  [12, 42, 25] suggested the distinguishability of BSs and BHs as well as NSs based on tidal deformability. Krishnendu et al. [28, 29] showed that the spin-induced higher multipole moments can be used to distinguish BH from other ECOs/BH mimickers. Later on, Pacilio et al. [34, 48] showed that coherently including the spin-induced quadrupole moment in the waveform along with tidal effects can be used to obtain complementary bounds on bosonic self-interactions with next-generation GW detectors. All these studies are restricted to binary systems and the inspiral phase of mergers. GWs from isolated BSs are again restricted to accreting cases. In this work, we propose a new signature based on $f$-mode oscillations to delineate massive BSs in the strong-interaction limit from BHs and NSs.
To do this, we compute the non-radial quadrupolar fundamental frequency that couples the strongest with gravitational waves. The study of quasinormal modes (QNMs) is crucial as it carries information about the nature of the underlying compact object [10, 11]. There have been a handful of studies on BS QNMs [26, 52, 6, 32][50] is the first work studying $f$-modes for massive BSs in the strong-interaction limit. The study was restricted to NS mass BSs, and the parameters were restricted by the bounds on self-interaction cross section $100.1<\sigma/m<$ 10 [cm2/g]. Another work [14] appeared during the completion of this work that explored select BS models in the same regime.
In this letter, we show for the first time that scaling relations are followed for the $f$-mode frequency and damping time in the strong-interaction limit. This scaling result is significant, as it allows commenting on the $f$-mode characteristics of BSs with arbitrary self-interacting scalar DM parameters within the strong interaction limit. Using this, we make a detailed comparison of BS $l=2$ QNMs with those of BHs and NSs to comment on the distinguishability. The analysis and implications of the scaling, along with the detectability, will be reported in a separate publication [43].
Throughout this work, we use the convention $\hbar=c=1$ and $G=1/M_{Pl}^{2}$, where $M_{Pl}=1.22\times 10^{22}\text{MeV}=2.18\times 10^{-8}\text{kg}$ is the Planck mass.
II Model
For pure Bosonic dark stars, we describe DM by scalar field ($\phi$) self-interacting via quartic self-interactions as described in [17, 50]. The Lagrangian is given by
$\mathcal{L}=\frac{1}{2}\partial_{\mu}\phi^{*}\partial^{\mu}\phi+V(\phi)~{},$ (1)
$V(|\phi|)=\frac{1}{2}m^{2}|\phi|^{2}+\frac{1}{4}\lambda|\phi|^{4}~{},$ (2)
where $\lambda$ is the self-interaction parameter and $m$ is the DM particle mass. The self-interaction is negligible only if $\lambda|\phi|^{4}/m^{2}|\phi|^{2}\ll 1$ [17], where $|\phi|\sim M_{Pl}$. A dimensionless parameter is defined as
$\Lambda\equiv\frac{\lambda M_{Pl}^{2}}{4\pi m^{2}}~{}.$ (3)
For $\Lambda\gg 1$[17] showed that massive stars can be formed with mass resembling that of a degenerate fermionic star. In this case, $M_{max}=0.22\sqrt{\Lambda}M_{Pl}^{2}/m=0.06\sqrt{\lambda}M_{pl}^{3}/m^{2}$ [17]. Moreover, $\Lambda\gg 1$, can be achieved for $\lambda\ll 1$. In this limit, even if $\lambda\ll 1$, the interaction is important, and the macroscopic properties are dictated by the strength ($\sqrt{\lambda}/m^{2}$).
We consider here the case of self-interacting massive BSs, for which $\Lambda\gg 1$, which is referred to as the strong-interaction limit [3]. In this limit, the Einstein-Klein-Gordon system of equations for the scalar field resembles that of a perfect fluid star with the effective equation of state given by [17, 38, 16],
$p=\frac{m^{4}}{9\lambda}\left(\sqrt{1+\frac{3\lambda}{m^{4}}\rho}-1\right)^{2}~{}.$ (4)
III Scaling
The mass of massive BSs is known to scale as $\sqrt{\lambda}M_{pl}^{3}/m^{2}$ [17] while its radius scales as $\sqrt{\lambda}M_{Pl}/m^{2}$. We define a quantity as done in [16]:
$x\equiv\sqrt{\lambda}/m^{2}~{}.$ (5)
[17] showed that the energy density and pressure scale as $\rho^{\prime}=\rho x^{2},~{}p^{\prime}=px^{2}$ to obtain the scaled version of the equation of state
$p^{\prime}=\frac{1}{9}(\sqrt{1+3\rho^{\prime}}-1)^{2}~{}.$ (6)
The dimensionless mass ($M^{\prime}$) and radius ($R^{\prime}$) can then be defined for BSs as $M^{\prime}=M/(xM_{Pl}^{3})~{},~{}R^{\prime}=R/(xM_{Pl})$ and can be used to obtain scaled TOV equations.
Here, we are interested in the $f$-mode frequencies and damping times. We introduce the scaling for the complex $f$-mode eigenfrequency ($\omega$) for the first time which is given by
$\boxed{\omega=\omega^{\prime}/(xM_{Pl})}~{}.$ (7)
This means the $f$-mode frequency ($f$) and damping time $\tau$ scale as
$f$ $=f^{\prime}/(xM_{Pl})~{},$ (8)
$\tau$ $=\tau^{\prime}(xM_{Pl})~{}.$ (9)
We explicitly derive this scaling using the $f$-mode perturbation equations for the case of Cowling approximation (where the background metric is held constant) as well as in full general relativity (see supplementary material).
IV $f$-modes
In this section, we discuss the $f$-mode characteristics of massive BSs in the strong-interaction limit. When the scaling as mentioned in Eq. 7 is used, the $f$-mode equations become completely independent of the parameter $x$ containing the DM parameters (see supplementary material).
[…]
In Fig. 1, we first plot dimensionless $f^{\prime}$ as a function of $M^{\prime}$ and compactness $C=M/R=M^{\prime}/R^{\prime}$. We get unique curves from primed quantities. This is because the $f$-mode equations and perturbation equations are completely independent of model parameters when written in scaled coordinates. We can then use scaling $f^{\prime}=f(xM_{Pl})$, $M^{\prime}=M/(xM_{Pl}^{3})$ and $C^{\prime}=C$ to obtain the curves for any set of $(\lambda,m)$ parameters.
Figure 1: The scaled $f$-mode frequency ($f^{\prime}=f(xM_{Pl})$) for massive BSs in the strong interaction limit ($\Lambda\gg 1$) as a function of scaled mass $M^{\prime}=M/(xM_{Pl}^{3})$ and compactness $C$.
Figure 2: The scaled $f$-mode damping time ($\tau^{\prime}=\tau/(xM_{Pl})$) for massive BSs in the strong interaction limit ($\Lambda\gg 1$) as a function of scaled mass $M^{\prime}=M/(xM_{Pl}^{3})$ and compactness $C$.
$f^{\prime}$ is an increasing function of $M^{\prime}$ and $C$ and reaches a maximum value of
$\boxed{f^{\prime}_{max}=0.21}$ (10)
for the maximum mass configuration of BS. Thus, for arbitrary DM parameters, the maximum possible BS $f$-mode frequency corresponding to the maximum compactness configuration is given by $f_{max}=0.21m^{2}/(\sqrt{\lambda}M_{Pl})$ (see Eq. 8). The increasing behavior is expected as the frequency is known to increase with the average density ($f\sim\sqrt{M/R^{3}}$), and in the case of BSs, it is known that mass increases while the radius decreases as we increase the central density.
Next, we show $\tau^{\prime}$ as a function of $M^{\prime}$ and $C^{\prime}$ in Fig. 2. $\tau^{\prime}$ is decreasing function of $M^{\prime}$ and $C^{\prime}$. We use the log scale for $\tau^{\prime}$ as it varies by orders of magnitude. Again, we find that scaled relations are independent of model parameters, i.e., for a given $M^{\prime}$ (or $C^{\prime}$), we have a unique $\tau^{\prime}$. The minimum value of $\tau^{\prime}$ corresponding to the maximum mass configuration is
$\boxed{\tau^{\prime}_{min}=1900}~{}.$ (11)
This single solution is sufficient to reproduce all the results and for the entire parameter space in the strong coupling regime ($\Lambda\gg 1$) using the relations $\tau^{\prime}=\tau/(xM_{Pl})$, $M^{\prime}=M/(xM_{Pl}^{3})$ and $C^{\prime}=C$.
Combining these two results, we find that $N=f\tau=f^{\prime}\tau^{\prime}$ is a decreasing function of mass and compactness. This is expected as $\tau$ falls on a logarithmic scale. $N$ represents the number of oscillations that the BS undergoes before the amplitude of the oscillations damps by a factor of $e$. Thus, corresponding to the maximum compactness configuration, we get a minimum of
$\boxed{N_{min}=406}$ (12)
oscillations. Thus, BSs with any microscopic model parameters and compactness would undergo at least $N_{min}$ number of oscillations if $f$-modes are excited.
We have performed a detailed analysis of these solutions and also derived analytical fits for these quantities. This will be reported in a separate publication [43].
V Distinguishability/ Comparison with neutron stars and black holes
It would be interesting to compare the BS frequencies with those of the QNMs of other astrophysical compact objects. Here, we compare the fundamental frequency for NSs and BHs, which are dominant modes in the context of GW.
For the case of BSs, the $f$-mode frequency as a function of mass can be derived using the scaling relations. We have $f=f^{\prime}(C)/(xM_{Pl})$. We can then use the scaling of mass to eliminate $x$. Thus, we get a relation for BS $f$-mode solely as a function of mass and compactness.
$f_{BS}=f^{\prime}(C)M^{\prime}(C)\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)$ (13)
$\boxed{f_{BS}=202.67f^{\prime}(C)M^{\prime}(C)\left(\frac{M_{\odot}}{M_{BS}}\right)\text{kHz}}~{}.$ (14)
This is a very useful relation as the frequency can be computed using just two physical parameters, i.e., $M$ and $C$. We can use the unique solutions obtained in the previous section to evaluate $f^{\prime}(C)$ and $M^{\prime}(C)$ (see Fig. 1). The analytical fits for these solutions will be provided in [43]. For both cases, the fits are accurate within $0.25\%$ for $M^{\prime}(C)$ and $1\%$ for $f^{\prime}(C)$. Thus, the error in Eq. 13 is less than $1.25\%$ if one uses the proposed analytical relations. The blue regions in Fig. 3 are obtained by varying the compactness for a given mass. The case of maximum compactness (dark blue curve in Fig. 3) is given by
$f_{BS,max}$ $=f^{\prime}(C_{max})M^{\prime}(C_{max})\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)$ (15)
$\approx 0.013\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)~{}.$ (16)
Thus we arrive at the equation,
$\boxed{f_{BS,max}=2.67\left(\frac{M_{\odot}}{M_{BS}}\right)\text{kHz}}~{}.$ (17)
This only depends on the mass of the BS, as in the case of BHs.
Figure 3: $f$-mode frequency as a function of mass for (a) NS, (b) stellar-mass BH, and (c) SMBH mass range for massive BSs. The black line is the corresponding BH QNM frequency [27]
The $f$-mode oscillations of NSs have been studied in great detail. Given the underlying uncertainty in NS EoS, we obtain a band in the $f$-$M$ space. NSs lie in the mass range of 1-2.5 $M_{\odot}$ with compactness in the range 0.1-0.3. We show this mass range in Fig. 3(a). The blue contour shows the region covered by BSs. The upper boundary corresponds to $C=C_{max}=0.16$. Silver curves show the $f-M$ region spanned by NSs due to the uncertainty in the microscopic equation of state (taken from [37]). We see a partial overlap with the NS band, but BS frequencies in the higher mass range are lower than that of NSs. Thus, GW observation of $f$-mode oscillations of a compact object in the NS mass range but outside the expected NS frequency band would be a signature for BS.
Astrophysical BHs are not expected in the mass range of NSs. However, PBHs, if they exist, may have mass in the solar and sub-solar mass range. On the other hand, stellar mass BHs fall in the range 5-100 $M_{\odot}$ and supermassive BHs (SMBH) in galactic centers have masses $10^{6}$-$10^{9}$ $M_{\odot}$. We show these two mass ranges Figs. 3(b) and 3(c), respectively. BHs with masses larger than 100 $M_{\odot}$ are possible via mergers of lighter stellar mass BHs [2]. We show the $n=0$, $l=2$ fundamental $f$-mode QNM frequency [27] by the black curves in Fig. 3. This is given by $\omega M=0.37+0.19a$ [27] in geometrical units ($G=c=1$), $a$ is the angular momentum, which after reverting to natural units for the non-spinning case means
$f_{BH}$ $=\left(\frac{0.37}{2\pi}\right)\times\frac{1}{GM_{BH}}$
$\approx 0.059\left(\frac{M_{Pl}^{2}}{M_{BH}}\right)$ (18)
$=12.05\left(\frac{M_{\odot}}{M_{BH}}\right)\text{kHz}~{}.$ (19)
The BH frequency only depends on the mass of a non-rotating, chargeless BH [27]. The BH frequency is also consistently higher than that of BS by a certain factor. This is expected from the high compactness of BH ($C=0.5$), which is higher by a factor of $3.125$ compared to the most compact BS. This factor is readily calculable using Eqs. 16 and V. We define a ratio ($\mathcal{R_{BS/BH}}$) as the ratio of the fundamental QNM frequency of the most compact BS to that of BH of the same mass. This ratio is a constant, independent of mass, and is given by
$\boxed{\mathcal{R_{BS/BH}}=\frac{f_{BS,max}}{f_{BH}}=0.22}$ (20)
The BH frequency is higher throughout by a factor of $4.50.22^{-1}=$ 4.5. If we observe $f$-modes from dark compact objects consistently with a frequency less than $\mathcal{R_{BS/BH}}$ times the expected BH frequency, it would be a smoking gun for massive BSs. Note that a similar analysis for other potentials ($V(\phi)$), as well as for another type of BSs, e.g., Oscillatons and Proca Stars, is necessary to distinguish between different models for stars made of Bosonic DM. However, this still acts as a clear signature for distinguishing BHs from BSs.
In general, the ratio is given by
$\frac{f_{BS}}{f_{BH}}=16.98f^{\prime}(C)M^{\prime}(C)~{}.$ (21)
A similar relation was provided in [21] for the innermost stable circular orbit (ISCO) frequency ($f_{ISCO}$)
$f^{ISCO}=\frac{1}{2\pi}\left(\frac{C}{3}\right)^{3/2}\left(\frac{M_{Pl}^{2}}{M_{BH}}\right)~{}.$ (22)
Here, $M$ is the component mass in an equal mass binary. The ISCO frequency marks the end of the inspiral phase during the binary merger. For $C=0.5$, we get the ISCO frequency for BH. We compare the $f$-mode frequency with the ISCO as it marks the importance of dynamical tides during pre-merger in the case of a BS binary system. For BHs, the ratio is
$\frac{f_{BH}}{f_{BH}^{ISCO}}=0.37\times 6^{3/2}=5.44~{}.$ (23)
Note that this number is not as important as the one for BSs, as there are no tidal deformations in the case of binary inspiral BH. For BSs, we get
$\frac{f_{BS}}{f_{BS}^{ISCO}}=2\pi f^{\prime}(C)M^{\prime}(C)\left(\frac{3}{C}\right)^{3/2}~{}.$ (24)
For the case of BSs, the ratio depends on the compactness. The maximum value obtained for the most compact case is 6.79, which is higher than that for BH. The reason is that although the QNM frequency is higher for BH, the ISCO frequency is much lower in the case of BSs, which are much less compact. It is also a much higher ratio, especially when compared with NSs. The typical ISCO frequency is roughly between 500 Hz - 1kHz [7], and the $f$-mode frequencies lie between 1.5-2 kHz for typical NS masses. Hence, the ratios for NSs is roughly around the range 2-4. Therefore, we expect the dynamical $f$-mode effects to be less prominent in the case of BSs.
VI Summary
In this work, we focussed on Boson stars, a class of dark stars made of scalar bosons. These are modeled by a scalar field $\phi$ that interacts gravitationally and has self-interaction given by the potential $V(\phi)$. We focus on the simple case of quartic potential given by $V(\phi)=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4}$. The strong-interaction limit ($\Lambda=\lambda M_{Pl}^{2}/4\pi m^{2}/\gg 1$) leads to the formation of a “massive star” whose mass scales as $\propto\sqrt{\lambda}M_{Pl}^{3}/m^{2}$ resembling massive fermionic stars. Such stars are compact objects and can act as BH mimickers, forming a fraction of the DM in the universe.
We evaluate the $f$-mode characteristics for massive BSs within a full general-relativistic setup. We derive the $f$-mode frequency and damping time scaling relations and demonstrate for the first time that the scaled dimensionless $f$-mode characteristics ($f^{\prime}$ and $\tau^{\prime}$) are independent of the model parameters. We report the relations $f^{\prime}-M^{\prime}$, $f^{\prime}-C^{\prime}$, and $\tau^{\prime}-M^{\prime}$, $\tau^{\prime}-C^{\prime}$, which are unique curves for massive BSs. The scaling relations $f=f^{\prime}/(xM_{Pl})$ and $\tau=\tau^{\prime}(xM_{Pl})$ can be used to evaluate the $f$-mode characteristics for any parameter values in the available parameter space. $f^{\prime}$ is an increasing function of $M^{\prime}/C^{\prime}$ whereas $\tau^{\prime}$ is a decreasing one. Thus, the configuration corresponding to the maximum $M^{\prime}=0.06$ or $C^{\prime}=0.16$ was used to derive the maximum value of $f^{\prime}$ given by $f^{\prime}_{max}=0.21$ and a minimum value of $\tau^{\prime}$ given by $\tau^{\prime}_{min}=1900$. This also results in a minimum number of BS $f$-mode oscillations given by $N_{min}=406$.
We compare the BS $f$-mode frequencies with those of NSs and BH $l=2$ QNMs. For the maximum compactness configuration, the frequency is the highest and only depends on the mass, as in the case of BH. For such configurations, the BS $f$-mode frequency is lower than the corresponding BH frequency by a constant factor given by $\mathcal{R_{BS/BH}}=0.22$. Thus, observing BH mimickers with $f$-mode frequency consistently lower than 0.22 times that of the expected BH frequency would be a clear signature of massive BS. We also compare the BS $f$-mode frequency with the corresponding ISCO frequency and find that the ratio of $f$-mode frequency to that of ISCO is larger for BS than NS and BH, given their low compactness. Thus, we expect the effect of the dynamical tides in binary systems of massive BSs to be less prominent.
Flores et al. [50] is the first work to explore non-radial fundamental modes for massive BSs in the strong-coupling limit, followed by Celato et al. [14]. These works explored BS configurations for select model parameters. The scalings with $x$ introduced in this work can be used to compute $f$-mode characteristics for any model parameters using a single solution, establishing that there is no need to study $f$-modes separately for different parameter values.
A list of observational signatures for BSs and a comparison of $f_{ISCO}$ between ECOs and BHs was provided in [21]. This work extends the list to include the fundamental QNM frequencies and demonstrates that $f$-modes can be used as a tool to distinguish BSs from NSs and BHs.
Future: The results presented here can be used for searches of massive BS in the strong-interaction limit where the effect of $f$-modes is important. These would be relevant in the future when $f$-mode oscillations become detectable. We will discuss the detectability of $f$-modes from BSs in the other work [43].
In this work, we have compared BS $f$-modes with those of NSs and BHs. This can be further extended to include NSs with exotic matter such as hybrid stars, strange stars, DM admixed NSs, etc. Here we considered scalar DM with $\phi^{4}$ potential resulting in $C_{max}\approx 0.16$. Higher $C_{max}$ can be achieved with $\phi^{n}$ potential for $n>4$ [35] possibly resulting in a higher $f^{\prime}_{max}$ and a higher $\mathcal{R}$. This can lead to configurations filling the gap between NSs and BHs (see Fig. 3), making it distinguishable from NSs as well. A complete analysis of ECOs would also require consideration of other scalar DM potentials, vector DM, and other ECOs such as fermion stars, gravastars, wormholes, fuzzballs, etc. [13]. The work can also be extended to include spinning BSs.
Acknowledgements.
L.S. and J.S.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 ‘Stronginteraction matter under extreme conditions” – project no. 315477589 – TRR 211.
Supplementary material
Appendix A Scaling of $f$-mode equations: Cowling Formalism
We will now see how the $f$-mode equations and the $f$-mode parameters get modified when scaling and scaled EoS are applied. We first look at the equations in Cowling approximation where metric perturbations are neglected, assuming a weak gravitational field. We do not use the Cowling formalism in this work as we solve the equations in full general relativity. However, we demonstrate that the scaling holds for solutions in Cowling approximation as well for the sake of completeness and the benefit of any future works.
The spacetime metric for a spherically symmetric background is given by
$ds^{2}=-e^{2\Phi(r)}dt^{2}+e^{2\Lambda(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}{\theta}d\phi^{2}$ (25)
Here,
$\frac{d\Phi(r)}{dr}$ $=\frac{-1}{\epsilon(r)+p(r)}\frac{dp}{dr}$ (26)
$\frac{d\Phi^{\prime}(r^{\prime})}{dr^{\prime}}$ $=\frac{-1}{\epsilon^{\prime}(r^{\prime})+p^{\prime}(r^{\prime})}\frac{dp^{\prime}}{dr^{\prime}}$ (27)
Thus, $\Phi^{\prime}=\Phi$. Similarly, $\Lambda^{\prime}=\Lambda$. This metric function $\Lambda(r)$ is different from the dimensionless tidal deformability. The differential equations to solve for the $f$-mode frequencies are given by [45, 49, 40, 36]:
$\frac{dW(r)}{dr}$ $=\frac{d\epsilon}{dp}\left[\omega^{2}r^{2}e^{\Lambda(r)-2\phi(r)}V(r)+\frac{d\Phi(r)}{dr}W(r)\right]$
$-l(l+1)e^{\Lambda(r)}V(r)$ (28)
$\frac{dV(r)}{dr}$ $=2\frac{d\Phi(r)}{dr}V(r)-\frac{1}{r^{2}}e^{\Lambda(r)}W(r)$ (29)
The functions $V(r)$ and $W(r)$ are associated with the Lagrange displacement vector of the perturbed fluid. Now we write these in terms of our new coordinates, $r=r^{\prime}xM_{Pl}$. We see the $\omega$ (the $f$-mode frequency) appears as the product ($\omega r$). So we scale $\omega$ as inverse of $r$, $\omega=\omega^{\prime}/(xM_{Pl})$ such that $\omega r=\omega^{\prime}r$. The equation becomes:
$\frac{dW(r)}{dr^{\prime}xM_{Pl}}$ $=\frac{d\epsilon^{\prime}}{dp^{\prime}}\left[\omega^{\prime 2}r^{\prime 2}e^{\Lambda^{\prime}(r^{\prime})-2\phi^{\prime}(r^{\prime})}V(r)+\frac{d\Phi^{\prime}(r^{\prime})}{dr^{\prime}xM_{Pl}}W(r)\right]-l(l+1)e^{\Lambda^{\prime}(r)}V(r)$ (30)
$\frac{dV(r)}{dr^{\prime}xM_{Pl}}$ $=2\frac{d\Phi^{\prime}(r^{\prime})}{dr^{\prime}xM_{Pl}}V(r)-\frac{1}{r^{\prime 2}x^{2}M_{Pl}^{2}}e^{\Lambda^{\prime}(r^{\prime})}W(r)$ (31)
The solutions at the center of the star are given by the following ansatz
$W(r)=Ar^{l+1},$ $\>V(r)=-\frac{A}{l}r^{l}~{},$ (32)
$\implies W^{\prime}(r^{\prime})=Ar^{\prime l+1},$ $\>V^{\prime}(r^{\prime})=-\frac{A}{l}r^{\prime l}~{},$ (33)
where $A$ is the amplitude of oscillations. We see that $W$ scales as $r^{l+1}$ and $V$ as $r^{l}$. We use, $W(r)=W^{\prime}(r^{\prime})(xM_{Pl})^{l+1}$ and $V(r)=V^{\prime}(r^{\prime})(xM_{Pl})^{l}$. Adopting this scaling, we find the primed equations remain unchanged and are free of $x$. Thus, the solutions do not depend on model parameters. The desired eigenfrequency is obtained by imposing the boundary condition
$\omega^{2}e^{\Lambda(R)-2\Phi(R)}V(R)$ $+\frac{1}{R^{2}}\frac{d\Phi(r)}{dr}\Big{|}_{r=R}W(R)=0~{},$ (34)
$\implies\omega^{\prime 2}e^{\Lambda^{\prime}(R^{\prime})-2\Phi^{\prime}(R^{\prime})}V^{\prime}(R^{\prime})$ $+\frac{1}{R^{\prime 2}}\frac{d\Phi^{\prime}(r)^{\prime}}{dr^{\prime}}\Big{|}_{r=R}W^{\prime}(R^{\prime})=0~{},$ (35)
which comes from vanishing perturbed Lagrangian pressure at the surface. Thus, we can solve the eigenfrequencies $\omega^{\prime}$ in primed coordinates (EoS independent) for different configurations $M^{\prime}$ or $R^{\prime}$ and use the $\omega=\omega^{\prime}/(xM_{Pl})$ to get the $f$-mode frequencies for BSs with arbitrary $(\lambda,m)$.
Appendix B $f$-mode equations: Full General Relativity
Here, we present the equations used to solve the complex QNM frequencies using full-general relativity, where metric perturbations are coupled to matter. Then we provide the scaling of various quantities that make the equations independent of the model parameters.
B.1 Perturbations Inside the Star
The perturbed metric ($ds^{2}_{p}$) can be written as [46],
$ds^{2}_{p}=ds^{2}+h_{\mu\nu}dx^{\mu}dx^{\nu}~{}.$ (36)
For the even-parity (polar) perturbations where the GW and matter perturbations are coupled, the $h_{\mu\nu}$ can be expressed as [44, 46],
$h_{\mu\nu}=\begin{pmatrix}r^{l}He^{2\Phi}&i\omega r^{l+1}H_{1}&0&0\\ i\omega r^{l+1}H_{1}&r^{l}He^{2\lambda}&0&0\\ 0&0&r^{l+2}K&0\\ 0&0&0&r^{l+2}Ksin^{2}{\theta}\end{pmatrix}Y^{l}_{m}e^{i\omega t}~{},$
where $Y_{m}^{l}$ are spherical harmonics. $H,\ H_{1},\ K$ are perturbed metric functions and are functions of the radial coordinate $r$.
[…]
$\zeta^{r}=\frac{r^{l}}{r}e^{-\lambda}W(r)Y^{l}_{m}e^{i\omega t}~{};\quad\zeta^{\theta}=\frac{-r^{l}}{r^{2}}V(r)\frac{\partial Y^{l}_{m}}{\partial\theta}e^{i\omega t}~{};\quad\zeta^{\phi}=\frac{-r^{l}}{r^{2}sin^{2}\theta}V(r)\frac{\partial Y^{l}_{m}}{\partial\phi}e^{i\omega t}~{},$ (38)
where $W,V$ are amplitudes of the radial and transverse fluid perturbations. The equations governing these perturbation functions and the metric perturbations inside the star are given by [44],
$\frac{dH_{1}}{dr}$ $=$ $\frac{-1}{r}\left[l+1+\frac{2m}{r}e^{2\lambda}+4\pi r^{2}e^{2\lambda}G\left(p-\epsilon\right)\right]H_{1}+\frac{1}{r}e^{2\lambda}\left[H+K+16\pi G\left(p+\epsilon\right)V\right]\ ,$ (39)
$\frac{dK}{dr}$ $=$ $\frac{l(l+1)}{2r}H_{1}+\frac{1}{r}H-\left(\frac{l+1}{r}-\frac{d\Phi}{dr}\right)K+\frac{8\pi}{r}G\left(p+\epsilon\right)e^{\lambda}W\ ,$ (40)
$\frac{dW}{dr}$ $=$ $re^{\lambda}\left[\frac{1}{\gamma Gp}e^{-\Phi}X-\frac{l(l+1)}{r^{2}}V-\frac{1}{2}H-K\right]-\frac{l+1}{r}W\ ,$ (41)
$\frac{dX}{dr}$ $=$ $\frac{-l}{r}X+G(p+\epsilon)e^{\Phi}\Bigg{[}\frac{1}{2}\left(\frac{d\Phi}{dr}-\frac{1}{r}\right)H-\frac{1}{2}\left(\omega^{2}re^{-2\Phi}+\frac{l(l+1)}{2r}\right)H_{1}$ (42)
$+$ $\left(\frac{1}{2r}-\frac{3}{2}\frac{d\Phi}{dr}\right)K-\frac{1}{r}\left[\omega^{2}\frac{e^{\lambda}}{e^{2\Phi}}+4\pi G(p+\epsilon)e^{\lambda}-r^{2}\frac{d}{dr}\left(\frac{e^{-\lambda}}{r^{2}}\frac{d\Phi}{dr}\right)\right]W$
-- $\frac{l(l+1)}{r^{2}}\frac{d\Phi}{dr}V\Bigg{]}\ ,$
$\left[1-\frac{3m}{r}-\frac{l(l+1)}{2}-4\pi r^{2}Gp\right]H-8\pi r^{2}e^{-\Phi}X+r^{2}e^{-2\lambda}\left[\omega^{2}e^{-2\Phi}-\frac{l(l+1)}{2r}\frac{d\Phi}{dr}\right]H_{1}$ (44)
-- $\left[1+\omega^{2}r^{2}e^{-2\Phi}-\frac{l(l+1)}{2}-(r-3m-4\pi r^{3}Gp)\frac{d\Phi}{dr}\right]K=0$
$e^{2\Phi}\left[e^{-\phi}X+\frac{e^{-\lambda}}{r}G\frac{dp}{dr}W+\frac{(p+\epsilon)}{2}H\right]-\omega^{2}G(p+\epsilon)V=0~{},$
where $X$ is introduced as [31, 44]
$X$ $=$ $\omega^{2}G(p+\epsilon)e^{-\Phi}V-\frac{We^{\Phi-\lambda}}{r}G\frac{dp}{dr}-\frac{1}{2}G(p+\epsilon)e^{\Phi}H\,,$
$m=m(r)$ is the enclosed mass of the star and $\gamma$ is the adiabatic index defined as
$\gamma=\frac{(p+\epsilon)}{p}\left(\frac{\partial p}{\partial\epsilon}\right)\bigg{|}_{ad}~{}.$ (46)
While solving these equations, we have to impose boundary conditions, i.e., the perturbation functions are finite throughout the interior of the star (particularly at the center, i.e., at $r=0$), and the perturbed pressure ($\Delta p$) vanishes at the surface. Function values at the center of the star can be found using the Taylor series expansion method described in Appendix B of [31] (see also Appendix A of [44]). The vanishing perturbed pressure at the stellar surface is equivalent to the condition $X(R)=0$ (as, $\Delta p=-r^{l}e^{-\Phi}X$). We follow the procedure described in [31] to find the unique solution for a given value of $l$ and $\omega$ satisfying all the boundary conditions inside the star.
B.2 Perturbations outside the star and complex eigenfrequencies
The perturbations outside the star are described by the Zerilli equation [53].
$\frac{d^{2}Z}{dr_{*}^{2}}+\omega^{2}Z=V_{Z}Z$ (47)
where $r_{*}=r+2m\log\left({\frac{r}{2m}-1}\right)$ is the tortoise co-ordinate and $V_{Z}$ is defined as  [53],
$V_{Z}$ $=$ $\frac{2(r-2m)}{r^{4}(nr+3m)^{2}}\Big{[}n^{2}(n+1)r^{3}+3n^{2}mr^{3}+9nm^{2}r+9m^{3}\Big{]}~{},$ (48)
where $n=\frac{1}{2}(l+2)(l-1)$. Asymptotically the wave solution to (47) can be expressed as (49),
$Z=A(\omega)Z_{in}$ $+$ $B(\omega)Z_{out}\,,$ (49)
$Z_{out}=e^{-i\omega r^{*}}\sum_{j=0}^{j=\infty}\alpha_{j}r^{-j}$ , $Z_{in}=e^{i\omega r^{*}}\sum_{j=0}^{j=\infty}\bar{\alpha}_{j}r^{-j}~{}.$
Keeping terms up to $j=2$ one finds,
$\alpha_{1}=-\frac{i}{\omega}(n+1)\alpha_{0},~{};\quad\alpha_{2}=\frac{-1}{2\omega^{2}}\left[n(n+1)-i3M\omega\left(1+\frac{2}{n}\right)\right]\alpha_{0}$ (50)
For initial boundary values of Zerilli functions, we use the method described in  [20, 18, 44]. Setting $m=M$ and perturbed fluid variables to 0 (i.e., $W=V=0$) outside the star, connection between the metric functions (37) with Zerilli function ($Z$ in Eq.(47)) can be written as,
$\begin{pmatrix}r^{l}K\\ r^{l+1}H_{1}\end{pmatrix}$ $=$ $\begin{pmatrix}\frac{n(n+1)r^{2}+3nMr+6M^{2}}{r^{2}(nr+3M)}&1\\ \frac{nr^{2}-3nMr-3M^{2}}{(r-2M)(nr+3M)}&\frac{r^{2}}{r-2M}\end{pmatrix}\begin{pmatrix}Z\\ \frac{dZ}{dr_{*}}\end{pmatrix}$ (51)
The initial boundary values of Zerilli functions are fixed using (51). The Zerilli equation (47) is then integrated numerically to infinity, and the complex coefficients $A(\omega),\ B(\omega)$ are obtained by matching the analytic expressions for $Z$ and $\frac{dZ}{dr_{*}}$ with the numerically obtained value of $Z$ and $\frac{dZ}{dr_{*}}$. The natural frequencies that are not driven by incoming gravitational radiation represent the quasi-normal mode frequencies. Mathematically, we find the complex roots of $A(\omega)=0$, representing the complex eigenfrequencies of QNMs.
B.3 Scaling
As the radius of the massive BS in the strong interaction limit scales as $R=R^{\prime}(xM_{Pl})$, we scale the radial coordinate as $r=r^{\prime}(xM_{Pl})$. The distance between two points also then becomes $ds=ds^{\prime}(xM_{Pl})$. Thus, the whole spatial scale is scaled up. $h_{\mu\nu}$, which is dimensionless, then remain unaltered under $r=r^{\prime}(xM_{Pl})$, $\theta^{\prime}=\theta$ and $\phi^{\prime}=\phi$. This is achieved if the quantities $H$ and $K$ scale as $H=H^{\prime}/(xM_{Pl})^{l}$ and $K=K^{\prime}/(xM_{Pl})^{l}$, $h_{\mu\nu}=h_{\mu\nu}^{\prime}$. Thus $H$ and $K$ scale as $1/r^{l}$. Using these and the perturbation equations, we arrive at the following transformations:
Scalings
$ds=ds^{\prime}(xM_{Pl})$ $H=H^{\prime}/(xM_{Pl})^{l}$ $m=m^{\prime}(xM_{Pl})$
$h_{\mu\nu}=h_{\mu\nu}^{\prime}$ $K=K^{\prime}/(xM_{Pl})^{l}$ $p=p^{\prime}/x^{2}$
$r=r^{\prime}(xM_{Pl})$ $H_{1}=H_{1}^{\prime}/(xM_{Pl})^{l}$ $\epsilon=\epsilon^{\prime}/x^{2}$
$t=t^{\prime}(xM_{Pl})$ $Y^{l}_{m}=Y^{l\prime}_{m}$ $\gamma=\gamma^{\prime}$
$\theta^{\prime}=\theta$ $W=W^{\prime}/(xM_{Pl})^{l-2}$ $\omega=\omega^{\prime}/(xM_{Pl})$
$\phi^{\prime}=\phi$ $V=V^{\prime}/(xM_{Pl})^{l-2}$ $X=X^{\prime}/(xM_{Pl})^{l+2}$
$\lambda=\lambda^{\prime}$ $\zeta^{r}=\zeta^{r\prime}(xM_{Pl})$ $Z=Z^{\prime}(xM_{Pl})$
$\Phi=\Phi^{\prime}$ $\zeta^{\theta}=\zeta^{\theta\prime}$ $V_{Z}=V_{Z}^{\prime}$
- $\zeta^{\phi}=\zeta^{\phi\prime}$
Applying these transformations and using $G=1/M_{Pl}^{2}$ in all the above equations, we find that they can be written completely in terms of the primed quantities, and all the factors of $x$ cancel out. Initially, the $x$ dependence was via the equation of state $p$ and $\epsilon$. Now $p^{\prime}$ and $\epsilon^{\prime}$ do not depend on $x$. Thus, the equations are completely free of the model parameters. For each central density $\rho^{\prime}_{c}$ or equivalently mass $M^{\prime}$, we get a unique solution for $\omega^{\prime}$. The scaling relation $\omega=\omega^{\prime}/(xM_{Pl})$ can then be used to obtain the $f$-mode frequency and damping time for any BS with DM model parameters $(\lambda,m)$, i.e., $\omega(x,M)=\omega^{\prime}(M^{\prime}=M/(xM_{Pl}^{3}))$. $\omega\propto f+i/\tau$. Thus we casn write $f=f^{\prime}/(xM_{Pl})$ and $\tau=\tau^{\prime}(xM_{Pl})$.
It is well-known that the quadrupolar fundamental non-radial frequency scales linearly with the square root of the average density of star [4, 5] ($f\sim\sqrt{GM/R^{3}}$). From the scaling of $M$ and $R$ mentioned above, we can say that the scaling of $f$ mentioned above is consistent with this. Thus, $f$ scales as the inverse of $r$.
Furthermore, the damping time should follow (see, Misner, Thorne, and Wheeler (1973) chapter 36.2); $\tau\sim R^{4}/(GM)^{3}$, which is also consistent with the derived scaling for $\tau$. The damping time is defined to be the inverse of the imaginary part of the complex eigenfrequency, so it should scale as the inverse of $f$. Moreover, the mass-scaled damping time ($GM/\tau$) shows universality with stellar compactness $C=M/R$ [47]. Hence, the initial guess would be that $\tau$ can be scaled as $r$.